# Rapidly-rotating Population III stellar models as a source of primary nitrogen

### v1

@@ 4 The rapidly-rotating grid
From Table 4, looking at column 10, we see that the non-rotating models show no increase in the surface He abundance between the end of the core H-burning phase and that of the core He-burning phase. This comes from the fact that the non-rotating models show no outer convective envelope during the core He-burning phase. They burn their helium in the core when the star is on the blue side of the HR diagram and thus no important outer convective zone is present. Intermediate and outer convective envelopes appear (if any) only after the core He-burning phase. In contrast, rapidly-rotating models show surface He enrichment that is significantly increased with respect to the value reached at the end of the core H-burning phase. This is a consequence of the outer and intermediate convective zones that are already appearing during the core He-burning phase in those models.
@@
Comparisons of the lifetimes of the core H- and He-burning phase for different initial rotations are given in columns 3 and 7 of Table 4. The MS phase of the rapidly-rotating 9 $M_{\odot}$ model is longer by about 50% than the duration for the non-rotating one. Much smaller effects are obtained in higher masses. The increase is only 10% for the 120 $M_{\odot}$ model. Since the differences between the moderately-rotating models (see Murphy et al., 2021a) and the present rapidly-rotating ones are modest during the MS phase for both the duration and the evolution in the HRD, we expect that the results on ionisation fluxes based on the moderate rotation rate models (Murphy et al., 2021b) are not changed much if rapidly-rotating models are used instead.
@@ 5 Primary nitrogen production in rotating models
Figure 6: The chemical structure as a function of the Lagrangian mass coordinate in the interior of the Pop III grids for $9M_{\odot}\leqslant M_{\rm ini}\leqslant$ 60 $M_{\odot}$. The figures are at the half-way through the core He-burning phase ($Y_{c}=0.5$) and at the last model that has been computed in different evolutionary stages, HeC is between the core He- and C-burning phase, CNe is the stage between the core C- and the Ne-burning phase, and O is during the core O-burning phase. The first column of plots corresponds to the moderately-rotating models with 40% of the critical rotation, and the middle and last columns to the rapidly-rotating models with 70% of the critical rotation. Each row refers to a different initial mass, the first row corresponds to 9 $M_{\odot}$, the second to 20 $M_{\odot}$, and the third to 60 $M_{\odot}$. Each coloured curve represents a different element, as it is noted on the upper left plot. The light brown zones are the convective areas of each model. The moderately-rotating models have been taken from Murphy et al. (2021a). The solid curves correspond to the most common isotope of an element, while the dashed curves to the second most common isotope.
Figure 7: The abundance of 14N as a function of the Lagrangian mass coordinate during different stages of the core He-burning phase for the 20 $M_{\odot}$ model with different initial rotations, zero (dotted curves), moderate (dashed curves), and fast (solid curves). The blue curves left correspond to the beginning of the core He-burning phase ($Y_{c}=0.9$), the green curves (middle) to the half-way through it ($Y_{c}=0.5$), and the red curves (right) to the end of it ($Y_{c}=0$). The non-rotating and moderately-rotating models were taken from Murphy et al. (2021a).
Figure 8: The evolution of the total mass of 14N inside the rapidly-rotating stellar models during their evolution. For the 20 $M_{\odot}$ model (orange), the non-rotating and moderately-rotating models are also indicated with the dotted and dashed curves. On the x-axis the fraction of burning phases is given, where the core H-burning phase is for 0--1, the core He-burning phase is for 1--2, and the advanced burning phases are for 2--3. The non-rotating and moderately-rotating models were taken from Murphy et al. (2021a).
@@
The first column shows the structure at the middle of the core He-burning phase in models starting with a rotation equal to 40% of the critical limit. We see that in those models some carbon and oxygen are brought by rotational mixing into the H-burning shell inducing a small production of nitrogen (see e.g. the green curve bump at a mass around 4 $M_{\odot}$ in the 20 $M_{\odot}$ model). Some of this nitrogen can diffuse into the He-burning core and be transformed into 22Ne (see the dashed purple curve into the He-burning core, e.g. below the mass coordinate 2.5 $M_{\odot}$).
@@
The panels in the middle column show models similar to those of the panels of the first column but now considering an initial rotation of 70% of the critical one. Here we see that the bumps in nitrogen are much more prominent. More specifically, comparing the rapidly-rotating models, we see that nitrogen dominates in the outer regions as we move towards to the surface in the 20 $M_{\odot}$ model. This exceptional fraction of nitrogen can be attributed to the presence of the convective shell in this model, a feature absent in the 9 model and occurring a later evolutionary phase in the 60 $M_{\odot}$. We can also see that 13C (see the red dashed curve) is produced in the H-burning shell (let us recall that 12C is transformed partially into 13C by the CNO cycle). Interestingly also in the 20 $M_{\odot}$ model, a significant quantity of 19F appears in the He-burning core (this point will be discussed in a dedicated paper, Tsiatsiou et al. (in prep.)). Finally, 22Ne also reaches a high abundance in the He-core (see the purple dashed curve) as a result of the diffusion/mixing of nitrogen into the He-burning core.
@@
The panels on the right show the same models as those of the middle panel but at a more advanced evolutionary stage, indicated in the caption of Fig. 6. We see that in general, the abundance of nitrogen has increased in the outer layers of the star. This is because of the continuing build-up of this element in the H-burning shell due to the diffusion of carbon and oxygen into it. The appearance of convective zones can redistribute nitrogen over larger regions in the envelope. We can note the very high level of nitrogen in the outer layers of the 20 and 60 $M_{\odot}$ models.
@@
Figure 7 allows us to see more clearly the impact of the initial rotation. These three panels compare the distribution of nitrogen as a function of the Lagrangian mass coordinate in the 20 $M_{\odot}$ model with three different initial rotations at three different evolutionary stages of the core He-burning phase. We can note that the nitrogen is mostly built in the early phases of the core He-burning. Besides, in the rapidly-rotating models ($\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$), the increase is impressive. We see that even the central regions are enriched in nitrogen. This is a consequence of the very large intermediate convective zone that actually extends nearly to the centre (as can be seen in the lower panel of Fig. 4) and thus brings nitrogen in very central regions. When He-burning restarts in the core, a new convective core appears that maintains for a while at least a high level of nitrogen abundance.
@@
In Fig. 8, the solid curves show how the integrated mass of nitrogen inside a model varies when the star evolves for different rapidly-rotating models. The integrated mass at a given time $t$ is given by $M_{{}^{14}\text{N}}=\int_{0}^{M_{tot}(t)}{}^{14}X(M_{r},t)\,{\rm d}M_{r}$, where $M_{tot}(t)$ is the actual total mass of the star at the time $t$, ${}^{14}X(M_{r},t)$ the mass fraction of 14N at the Lagrangian coordinate $M_{r}$ at time $t$.
@@
We see that in the 9 $M_{\odot}$ model, the increase of $M_{{}^{14}\text{N}}$ occurs progressively during the whole core He-burning phase. For the higher initial mass models, the increase is more step-like. As just seen above, it occurs at the very beginning of the core He-burning phase in the 20 $M_{\odot}$ model and at later and later core He-burning stage in more and more massive stars. In these higher initial mass models, convection plays an important role in connecting regions rich in both H- and He-burning products.
@@
We have also plotted for the 20 $M_{\odot}$ model the evolution of $M_{{}^{14}\text{N}}$ for the non-rotating and moderately-rotating models (respectively the orange dotted and dashed curves). Interestingly, even the non-rotating model shows some step-like increase of $M_{{}^{14}\text{N}}$ during the core C-burning phase. The moderately-rotating model does not show such a feature but we cannot discard the fact that it may present one in a more advanced stage. We can note that in non-rotating models such connection between the He- and H-burning regions leading to significant primary nitrogen production seems to be restricted to a narrow mass domain (Ekström et al., 2010; Limongi & Chieffi, 2012), while rotational mixing not only boosts the primary nitrogen production in this mass domain but also extends the mass domain over which this process occurs across the whole mass range of stars considered here (see Fig.  9).
@@
Moreover, in Fig. 10 we plot the integrated abundance of nitrogen (green) and oxygen (blue) as a function of the Lagrangian mass coordinate for the 20 $M_{\odot}$ rapidly-rotating model at its last evolutionary stage (during the core O-burning phase). We show in the plot where the remnant mass ($M_{\rm rem}$ in grey dashed line) is located for the model, as well as the typical mass of a neutron star (NS) ($M_{\rm NS}$ in light grey dashed line). It is clear that the selection of the remnant mass can have a strong impact on the oxygen yields. Typically we see in Fig. 10 that the integrated mass of nitrogen, starting the integration in mass from the surface increases when considering still deeper layers until the lagrangian mass around 11 $M_{\odot}$ is reached. Below that mass, the nitrogen mass does not longer increase because below this mass, nitrogen is completely destroyed. Oxygen remains abundant in much deeper layers since O-burning only occurs in the very central parts of the star. As a consequence, the integrated mass of oxygen continues to increase when deeper layers are added to the mass ejected.
@@
Table 2: Stellar yields of nitrogen and oxygen for different Pop III and very metal-poor stellar models. The asterisk after the initial mass indicates that the model is still during the core He-burning phase. The Pop III models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=$ 0, 0.2, and 0.4 are from Murphy et al. (2021a) and the models with $Z=10^{-5}$ are from Sibony et al. (submitted).
@@
We use here the classical definition of the stellar yield of a given isotope. It is the quantity of that isotope newly synthesised and ejected by the star. Since in the present models, the mass lost in the course of the evolution is negligible, only the mass ejected at the time of the core collapse has to be accounted for. Thus, we only need to know the remnant mass to calculate the stellar yields. To obtain the final fate and remnant mass we use the results of Farmer et al. (2019); Patton & Sukhbold (2020), following the method outlined in Sibony et al. (2023) and in Nandal, Sibony & Tsiatsiou (2024) for different CO core masses. The stellar yields are obtained by integrating the abundances of an element above the remnant mass and by subtracting the initial quantity of the element, as indicated in Maeder (1992). Of course for Pop III stellar models, the mass of a metal isotope ejected is the stellar yield of that isotope.
@@
The results are presented in Table LABEL:table:yields_Patton for different metallicities and different initial rotations. The initial mass is given in column 1. In columns 2 and 3, the CO core and the remnant masses are given, respectively. The CO core mass corresponds to the mass inside the Lagrangian mass coordinate where the mass fraction of helium drops below $10^{-2}$ at the end of the core He-burning phase. We also provide the possible final fate of the stellar models in column 4. Lastly, the stellar yields of nitrogen and oxygen are given in columns 5 and 6, respectively.
@@
The models have been computed until the end of the core He-burning phase (some have been pushed further), so the yields that we obtain may still be affected by the more advanced stages of the evolution. However, the changes for the two isotopes considered here remain modest and thus the yields obtained here are a good approximation of the yields that would be obtained from pre-SN models. Also, the bulk of nitrogen and oxygen are located sufficiently away from the central conditions so that they will be affected very little by explosive nucleosynthesis.
@@
Figure 9 shows the stellar yields of nitrogen and oxygen. The upper panels show the yields for Pop III models with different initial rotations. The non-rotating and moderately-rotating models have been taken from Murphy et al. (2021a). We have to note that the stellar yields from Murphy et al. (2021a) are re-calculated for this work, since we have a different method of calculating the remnant mass. In Murphy et al. (2021a) the remnant mass is calculated according to Maeder (1992). The middle panels show the yields from the rapidly-rotating Pop III models together with the moderately-rotating metal-poor models ($Z=10^{-5}$) from Sibony et al. (submitted). The two lower panels show the yields for Pop III models with different initial rotations and from different authors. The non-rotating models have been taken from Limongi & Chieffi (2012) and Murphy et al. (2021a), and the moderately-rotating from Ekström et al. (2008). We shall discuss these two lower panels in Sect. 6.3.
@@
The production of 14N is boosted in the rapidly-rotating models as can be seen by comparing the solid green curve with the dashed blue (no rotation) and orange (moderate rotation) curves. The moderately-rotating models show smaller yields of nitrogen than the non-rotating ones in the mass range between about 10 and 87 $M_{\odot}$. As explained in Sect. A.3, this results from very different behavior of the H-burning shell in the two models. Actually in the non-rotating models, the structure show that the H- and the He-burning shell are much nearer from each other than in the rotating model. This favours mixing between these shells in the non-rotating model after the core He-burning phase.
@@
Regarding the comparison between the Pop III models and those with $Z=10^{-5}$, we see that the rapidly-rotating Pop III stars may contribute even more than the moderately-rotating $Z=10^{-5}$ models (see middle left panel). The $Z=10^{-5}$ models in the high mass range lose significant amounts of mass by stellar winds. As a numerical example the 85 $M_{\odot}$ model at $Z=10^{-5}$ loses about 30 $M_{\odot}$. This has a strong impact on the structure, notably making the H-shell to be less active, more distant from the He-burning region and thus making primary nitrogen production less efficient. In Fig. 10, we show how the yield of nitrogen would vary when considering different remnant masses for the 20 $M_{\odot}$ model. For instance, for the remnant mass given in Table LABEL:table:yields_Patton, the yield is given by the ordinate of the intersection between the nitrogen curve and the vertical dashed line labelled $M_{\rm rem}$ in Fig. 10. Considering lower values for the remnant mass would not change the nitrogen yield much. Considering larger values would on the contrary significantly impact the yield, even suppressing it completely if the whole final mass of the star were to remain locked into the stellar remnant black hole.
@@
The yields for the main isotope of oxygen show significant differences between the rapidly-rotating, non-rotating and moderately-rotating models. Most of these differences reflect a complex interaction between the effects of rotation on the size of the CO core and the process of computing the remnant mass. Changing the remnant mass may strongly impact the yields of oxygen. Looking at Fig. 10, we see that in case the rapidly-rotating 20 $M_{\odot}$ model produces a 12.7 $M_{\odot}$ black hole or a 1.6 $M_{\odot}$ neutron star, the yield of oxygen varies from a value of 0.2 $M_{\odot}$ up to a value of nearly 6 $M_{\odot}$. This high sensitivity of the oxygen yields on the remnant mass explains the very large difference between the yields of the non-rotating/moderately-rotating 60 $M_{\odot}$ models and the rapidly-rotating one. In the rapidly-rotating model, a smaller CO core mass is produced (14.4 $M_{\odot}$ instead of 22.6 $M_{\odot}$ in the non-rotating model). This is a result, as already explained above, of the extended intermediate convective zone attached to the H-burning shell when a CNO shell boost occurs. This has the consequence of reducing the He- and then the CO core mass and finally the remnant mass in the rapidly-rotating model. It has a dramatic impact on the oxygen yields passing from a value equal to $4.4\times 10^{-7}$ $M_{\odot}$ in the non-rotating model to a value equal to 0.91 $M_{\odot}$ in the rapidly-rotating one. This illustrates the great sensitivity of the oxygen yield on the final fate of a star, a point that was already discussed in detail by Maeder (1992). Given this and since the final fate of massive stars is still a topic associated with great uncertainties, we provide in our webpage [^7] https://www.unige.ch/sciences/astro/evolution/en/database/ the structure of our last computed models, so that other remnant masses can be considered to obtain the yields.
@@
In the middle right panel of Fig. 9, the oxygen yields for the $Z=10^{-5}$ models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ are compared with the Pop III oxygen yield from models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$. Some masses show very similar values, and the others significant differences, without any systematic effects. Again here the behaviour is very dependent on the way the remnant masses are computed.
@@
Figure 9: The stellar yields of nitrogen (left) and oxygen (right) for Pop III and very low-metallicity ($Z=10^{-5}$) stars, for different initial rotations and from different authors (Ekström et al., 2012; Limongi & Chieffi, 2012; Murphy et al., 2021a, Sibony et al. submitted). The ranges of initial equatorial velocities corresponding to the different coloured curves are indicated.
@@
Figure 10: Variation of the mass of nitrogen and oxygen in the outer layers of the star as a function of the lagrangian mass taken as the lower limit of the integration. This lower limit can be called the mass of the remnant. The case of the rapidly-rotating 20 $M_{\odot}$ model is shown. The vertical grey dashed line denotes the remnant mass as adopted in the present paper, while the horizontal grey dashed line denotes the nitrogen yield. The lighter grey dashed line denotes the typical mass of a neutron star.

### v2

@@ 4 The rapidly rotating grid
From Table 4, looking at column 10, we see that the non-rotating models show no increase in the surface He abundance between the end of the core H-burning phase and that of the core He-burning phase. This comes from the fact that the non-rotating models show no outer convective envelope during the core He-burning phase. They burn their helium in the core when the star is on the blue side of the HR diagram, and thus no important outer convective zone is present. Intermediate and outer convective envelopes appear (if any) only after the core He-burning phase. In contrast, rapidly rotating models show surface He-enrichment that is significantly increased with respect to the value reached at the end of the core H-burning phase. This is a consequence of the outer and intermediate convective zones that already appear during the core He-burning phase in those models.
@@
Comparisons of the lifetimes of the core H- and He-burning phase for different initial rotations are given in columns 3 and 7 of Table 4. The MS phase of the rapidly rotating 9 $M_{\odot}$ model is about 50% longer than the duration of the non-rotating one. Much smaller effects are obtained in higher masses. The increase is only 10% for the 120 $M_{\odot}$ model. Since the differences between the moderately rotating models (see Murphy et al., 2021a) and the present rapidly rotating ones are modest during the MS phase for both the duration and the evolution in the HRD, we do not expect the results for ionisation fluxes based on the moderate rotation rate models (Murphy et al., 2021b) to change much if rapidly rotating models are used instead.
@@ 5 Primary nitrogen production in rotating models
Figure 6: Chemical structure as a function of the Lagrangian mass coordinate in the interior of the Pop III grids for $9M_{\odot}\leqslant M_{\rm ini}\leqslant$ 60 $M_{\odot}$. The figures are at halfway through the core He-burning phase ($Y_{c}=0.5$) and at the last model that has been computed in different evolutionary stages. HeC is between the core He- and C-burning phase, CNe is the stage between the core C- and the Ne-burning phase, and O is during the core O-burning phase. The first column of the plots corresponds to the moderately rotating models with 40% of the critical rotation, and the middle and last columns correspond to the rapidly rotating models with 70% of the critical rotation. Each row refers to a different initial mass: the first row corresponds to 9 $M_{\odot}$, the second to 20 $M_{\odot}$, and the third to 60 $M_{\odot}$. Each coloured curve represents a different element, as is noted in the upper left plot. The light brown zones are the convective areas of each model. The moderately rotating models have been taken from Murphy et al. (2021a). The solid curves correspond to the most common isotope of an element, and the dashed curves to the second most common isotope.
Figure 7: Abundance of 14N as a function of the Lagrangian mass coordinate during different stages of the core He-burning phase for the 20 $M_{\odot}$ model with different initial rotations: zero (dotted curves), moderate (dashed curves), and fast (solid curves). The blue curves (left) correspond to the beginning of the core He-burning phase ($Y_{c}=0.9$), the green curves (middle) to the halfway point ($Y_{c}=0.5$), and the red curves (right) to the end of it ($Y_{c}=0$). The non-rotating and moderately rotating models were taken from Murphy et al. (2021a).
Figure 8: Evolution of the total mass of 14N inside the rapidly rotating stellar models during their evolution. For the 20 $M_{\odot}$ model (orange), the non-rotating and moderately rotating models are also indicated by dotted and dashed curves. On the x-axis the fraction of burning phases is given, where the core H-burning phase is for 0--1, the core He-burning phase is for 1--2, and the advanced burning phases are for 2--3. The non-rotating and moderately rotating models were taken from Murphy et al. (2021a).
@@
The first column shows the structure in the middle of the core He-burning phase in models starting with a rotation equal to 40% of the critical limit. We see that in those models some carbon and oxygen are brought by rotational mixing into the H-burning shell, inducing the production of a small amount of nitrogen (see e.g. the green curve bump at a mass of around 4 $M_{\odot}$ in the 20 $M_{\odot}$ model). Some of this nitrogen can diffuse into the He-burning core and be transformed into 22Ne (see the dashed purple curve into the He-burning core, e.g. below the mass coordinate 2.5 $M_{\odot}$).
@@
The panels in the middle column show models similar to those of the panels of the first column but now considering an initial rotation of 70% of the critical one. Here we see that the bumps in nitrogen are much more prominent. More specifically, comparing the rapidly rotating models, we see that nitrogen dominates in the outer regions as we move towards to the surface in the 20 $M_{\odot}$ model. This exceptional fraction of nitrogen can be attributed to the presence of the convective shell in this model, a feature absent in the 9 model and occurring in a later evolutionary phase in the 60 $M_{\odot}$. We can also see that 13C (see the dashed red curve) is produced in the H-burning shell (let us recall that 12C is transformed partially into 13C by the CNO cycle). Interestingly, in the 20 $M_{\odot}$ model a significant quantity of 19F also appears in the He-burning core (this point will be discussed in a dedicated paper by Tsiatsiou et al. (in prep.)). Finally, 22Ne also reaches a high abundance in the He-core (see the dashed purple curve) as a result of the diffusion or mixing of nitrogen into the He-burning core.
@@
The panels on the right show the same models as those of the middle panel but at a more advanced evolutionary stage, indicated in the caption of Fig. 6. We see that in general the abundance of nitrogen has increased in the outer layers of the star. This is because of the continuing build-up of this element in the H-burning shell due to the diffusion of carbon and oxygen into it. The appearance of convective zones can redistribute nitrogen over larger regions in the envelope. We note the very high level of nitrogen in the outer layers of the 20 and 60 $M_{\odot}$ models.
@@
Figure 7 allows us to see more clearly the impact of the initial rotation. These three panels compare the distribution of nitrogen as a function of the Lagrangian mass coordinate in the 20 $M_{\odot}$ model with three different initial rotations at three different evolutionary stages of the core He-burning phase. We note that the nitrogen mostly builds up in the early phases of the core He-burning. Besides, in the rapidly rotating models ($\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$), the increase is impressive. We see that even the central regions are enriched in nitrogen. This is a consequence of the very large intermediate convective zone that actually extends nearly to the centre (as can be seen in the lower panel of Fig. 4), and thus brings nitrogen into very central regions. When He-burning restarts in the core, a new convective core appears that maintains, for a while at least, a high level of nitrogen abundance.
@@
In Fig. 8, the solid curves show how the integrated mass of nitrogen inside a model varies when the star evolves for different rapidly rotating models. The integrated mass at a given time, $t$, is given by $M_{{}^{14}\text{N}}=\int_{0}^{M_{tot}(t)}{}^{14}X(M_{r},t)\,{\rm d}M_{r}$, where $M_{tot}(t)$ is the actual total mass of the star at the time, $t$, and ${}^{14}X(M_{r},t)$ is the mass fraction of 14N at the Lagrangian coordinate, $M_{r}$, at time $t$.
@@
We see that in the 9 $M_{\odot}$ model, the increase in $M_{{}^{14}\text{N}}$ occurs progressively throughout the whole core He-burning phase. For the models with a higher initial mass, the increase is more step-like. As was just seen above, it occurs at the very beginning of the core He-burning phase in the 20 $M_{\odot}$ model and at later and later core He-burning stages in more and more massive stars. In these models with a higher initial mass, convection plays an important role in connecting regions rich in both H- and He-burning products.
@@
We have also plotted for the 20 $M_{\odot}$ model the evolution of $M_{{}^{14}\text{N}}$ for the non-rotating and moderately rotating models (the dotted and dashed orange curves, respectively). Interestingly, even the non-rotating model shows a step-like increase in $M_{{}^{14}\text{N}}$ during the core C-burning phase. The moderately rotating model does not show such a feature but we cannot discard the fact that it may present one in a more advanced stage. We note that in non-rotating models such a connection between the He- and H-burning regions leading to significant primary nitrogen production seems to be restricted to a narrow mass domain (Ekström et al., 2010; Limongi & Chieffi, 2012), while rotational mixing not only boosts the primary nitrogen production in this mass domain but also extends the mass domain over which this process occurs across the whole mass range of stars considered here (see Fig.  9).
@@
Moreover, in Fig. 10 we plot the integrated abundance of nitrogen (green) and oxygen (blue) as a function of the Lagrangian mass coordinate for the 20 $M_{\odot}$ rapidly rotating model in its last evolutionary stage (during the core O-burning phase). We show in the plot where the remnant mass ($M_{\rm rem}$, dashed grey line) is located for the model, as well as the typical mass of a neutron star ($M_{\rm NS}$, dashed light grey line). It is clear that the selection of the remnant mass can have a strong impact on the oxygen yields. Typically, we see in Fig. 10 that for the integrated mass of nitrogen, starting the integration in mass from the surface increases when considering still deeper layers until the Lagrangian mass around 11 $M_{\odot}$ is reached. Below that mass, the nitrogen mass no longer increases because the nitrogen has been completely destroyed. Oxygen remains abundant in much deeper layers since O-burning only occurs in the very central parts of the star. As a consequence, the integrated mass of oxygen continues to increase when deeper layers are added to the mass ejected.
@@
Table 2: Stellar yields of nitrogen and oxygen for different Pop III and very metal-poor stellar models. The asterisk after the initial mass indicates that the model is still in the core He-burning phase. The Pop III models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=$ 0, 0.2, and 0.4 are from Murphy et al. (2021a) and the models with $Z=10^{-5}$ are from Sibony et al. (2024).
@@
We use here the classical definition of the stellar yield of a given isotope. It is the quantity of the isotope that has been newly synthesised and ejected by the star. Since in the present models the mass lost in the course of the evolution is negligible, only the mass ejected at the time of the core collapse has to be accounted for. Thus, we only need to know the remnant mass to calculate the stellar yields. To obtain the final fate and remnant mass we used the results of Farmer et al. (2019); Patton & Sukhbold (2020), following the method outlined in Sibony et al. (2023) and in Nandal, Sibony, & Tsiatsiou (2024) for different CO core masses. The stellar yields were obtained by integrating the abundances of an element above the remnant mass and by subtracting the initial quantity of the element, as was indicated in Maeder (1992). Of course for Pop III stellar models, the mass of an ejected metal isotope is the stellar yield of that isotope.
@@
The results are presented in Table LABEL:table:yields_Patton for different metallicities and different initial rotations. The initial mass is given in column 1. In columns 2 and 3, the CO-core and the remnant masses are given, respectively. The CO-core mass corresponds to the mass inside the Lagrangian mass coordinate where the mass fraction of helium drops below $10^{-2}$ at the end of the core He-burning phase. We also provide the possible final fate of the stellar models in column 4. Lastly, the stellar yields of nitrogen and oxygen are given in columns 5 and 6, respectively.
@@
The models were computed until the end of the core He-burning phase (some have been pushed further), so the yields that we obtain may still be affected by the more advanced stages of the evolution. However, the changes for the two isotopes considered here remain modest, and thus the yields obtained here are a good approximation of the yields that would be obtained from pre-SN models. Also, the bulk of nitrogen and oxygen is located sufficiently far away from the central conditions for them to be affected very little by explosive nucleosynthesis.
@@
Figure 9 shows the stellar yields of nitrogen and oxygen. The upper panels show the yields for Pop III models with different initial rotations. The non-rotating and moderately rotating models have been taken from Murphy et al. (2021a). We note that the stellar yields from Murphy et al. (2021a) were re-calculated for this work, since we have a different method of calculating the remnant mass. In Murphy et al. (2021a) the remnant mass was calculated according to Maeder (1992). The middle panels show the yields from the rapidly rotating Pop III models, together with the moderately rotating metal-poor models ($Z=10^{-5}$) from Sibony et al. (2024). The two lower panels show the yields for Pop III models with different initial rotations and from different authors. The non-rotating models have been taken from Limongi & Chieffi (2012) and Murphy et al. (2021a), and the moderately rotating from Ekström et al. (2008). We discuss these two lower panels in Sect. 6.3.
@@
The production of 14N is boosted in the rapidly rotating models, as can be seen by comparing the solid green curve with the dashed blue (no rotation) and orange (moderate rotation) curves. The moderately rotating models show smaller yields of nitrogen than the non-rotating ones in the mass range between about 10 and 87 $M_{\odot}$. As is explained in Sect. A.3, this results from very different behavior of the H-burning shell in the two models. Actually, in the non-rotating models, the structure shows that the H- and the He-burning shell are much nearer to each other than in the rotating model. This favours mixing between these shells in the non-rotating model after the core He-burning phase.
@@
Regarding the comparison between the Pop III models and those with $Z=10^{-5}$, we see that the rapidly rotating Pop III stars may contribute even more than the moderately rotating $Z=10^{-5}$ models (see middle left panel). The $Z=10^{-5}$ models in the high mass range lose significant amounts of mass by stellar winds. As a numerical example, the 85 $M_{\odot}$ model at $Z=10^{-5}$ loses about 30 $M_{\odot}$. This has a strong impact on the structure, notably making the H-shell less active and more distant from the He-burning region, and thus making primary nitrogen production less efficient. In Fig. 10, we show how the yield of nitrogen would vary when considering different remnant masses for the 20 $M_{\odot}$ model. For instance, for the remnant mass given in Table LABEL:table:yields_Patton, the yield is given by the co-ordinates of the intersection between the nitrogen curve and the vertical dashed line labelled $M_{\rm rem}$ in Fig. 10. Considering lower values for the remnant mass would not change the nitrogen yield much. Considering larger values would, on the contrary, significantly impact the yield, even suppressing it completely if the whole final mass of the star were to remain locked into the stellar remnant black hole.
@@
The yields for the main isotope of oxygen show significant differences between the rapidly rotating non-rotating, and moderately rotating models. Most of these differences reflect a complex interaction between the effects of rotation on the size of the CO-core and the process of computing the remnant mass. Changing the remnant mass may strongly impact the yields of oxygen. Looking at Fig. 10, we see that in the case in which the rapidly rotating 20 $M_{\odot}$ model produces a 12.7 $M_{\odot}$ black hole or a 1.6 $M_{\odot}$ NS, the yield of oxygen varies from a value of 0.2 $M_{\odot}$ up to a value of nearly 6 $M_{\odot}$. This high sensitivity of the oxygen yields on the remnant mass explains the very large difference between the yields of the non-rotating or moderately rotating 60 $M_{\odot}$ models and the rapidly rotating one. In the rapidly rotating model, a smaller CO-core mass is produced (14.4 $M_{\odot}$ instead of 22.6 $M_{\odot}$ in the non-rotating model). This is a result, as was already explained above, of the extended intermediate convective zone attached to the H-burning shell when a CNO shell boost occurs. This has the consequence of reducing the He- and then the CO-core mass and finally the remnant mass in the rapidly rotating model. It has a dramatic impact on the oxygen yields, passing from a value equal to $4.4\times 10^{-7}$ $M_{\odot}$ in the non-rotating model to a value equal to 0.91 $M_{\odot}$ in the rapidly rotating one. This illustrates the great sensitivity of the oxygen yield to the final fate of a star, a point that was already discussed in detail by Maeder (1992). Given this, and since the final fate of massive stars is still a topic associated with great uncertainties, we provide on our webpage [^7] https://www.unige.ch/sciences/astro/evolution/en/database/ the structure of our last computed models, so that other remnant masses can be considered to obtain the yields.
@@
In the middle right panel of Fig. 9, the oxygen yields for the $Z=10^{-5}$ models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.4$ are compared with the Pop III oxygen yield from models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$. Some masses show very similar values, and others significant differences, without any systematic effects. Again, here the behaviour is very dependent on the way the remnant masses were computed.
@@
Figure 9: Stellar yields of nitrogen (left) and oxygen (right) for Pop III and very low-metallicity ($Z=10^{-5}$) stars, for different initial rotations and from different authors (Ekström et al., 2012; Limongi & Chieffi, 2012; Murphy et al., 2021a; Sibony et al., 2024). The ranges of initial equatorial velocities corresponding to the different coloured curves are indicated.
@@
Figure 10: Variation in the mass of nitrogen and oxygen in the outer layers of the star as a function of the Lagrangian mass taken as the lower limit of the integration. This lower limit can be called the mass of the remnant. The case of the rapidly rotating 20 $M_{\odot}$ model is shown. The vertical dashed grey line denotes the remnant mass adopted in the present paper, while the horizontal dashed grey line denotes the nitrogen yield. The dashed lighter grey line denotes the typical mass of an NS.